# Distribution System Reconfiguration to Mitigate Load Altering Attacks via Stackelberg Games

### v5

@@ VI Conclusions
In this paper, we investigate the impact of LAAs on DNs. We derive a set of closed-form expressions for the power flow of DNs to determine bus voltages in the presence of voltage-dependent loads, with or without LAA. Then, we introduce a sequential game-theoretic approach to mitigate LAAs in DNs by network reconfiguration with minimum possible switching. Furthermore, we take into account the uncertainties in the attack localization by introducing a probability distribution over the potentially attacked nodes. To enhance the sustainability and computational speed of the Stackelberg equilibrium, a Bayesian optimization algorithm was implemented to reduce the computational burden. The proposed mitigation scheme is capable of keeping the voltage within the desired constraints. Building on our mitigation method, future work will investigate the cyber recovery step, which includes isolating the attack. Other interesting extensions are: dynamic LAAs, integrating distributed energy resources in the DN model, including the transient dynamics of the grid’s voltage, etc.

### v6

@@ VI Conclusions and Future Works
In this paper, we investigate the impact of LAAs on DNs. We derive a set of closed-form expressions for the power flow of DNs to determine bus voltages in the presence of voltage-dependent loads, with or without LAA. Then, we introduce a sequential game-theoretic approach to mitigate LAAs in DNs by network reconfiguration with minimum possible switching and utilizing the flexibility of inverter-based DERs outputs. Furthermore, we take into account the uncertainties in the attack localization by introducing a probability distribution over the potentially attacked nodes. To enhance the sustainability and computational speed of the Stackelberg equilibrium, a Bayesian optimization algorithm was implemented to reduce the computational burden. The proposed mitigation scheme is capable of keeping the voltage within the desired constraints. Building on our mitigation method, future work will investigate the cyber recovery step, which includes isolating the attack. Other interesting extensions include dynamic LAAs and incorporating the transient dynamics of the grid’s voltage, among others. Furthermore, optimal auxiliary lines installation is another potential addition to this work, which should be pursued in the future.
Our study provides a foundational one-shot attack-defense model. This work can be extended by future research to develop a multi-stage game-theoretic formulation for analyzing more sophisticated interactions.